The increase or decrease in the price of a stock between the beginning and the end of a trading day is assumed to be an equally likely random event. What is the probability that a stock will show a decrease in its closing price on three consecutive days?
The probability that a stock will show a decrease in its closing price on three consecutive days is?

Answer by stanbon(75887)

The increase or decrease in the price of a stock between the beginning and the end of a trading day is assumed to be an equally likely random event.
P(increase) = P(decrease) = 1/2
----------------------
What is the probability that a stock will show a decrease in its closing price on three consecutive days?
-----
Answer:: (1/2)^3 = 1/8
===========================
Cheers,
Stan H.
===============